# Topological phase transitions in perovskite superlattices driven by temperature, electric field, and doping

### v2

@@ Abstract
Many dipolar topological structures have been experimentally demonstrated in (PbTiO3)n/(SrTiO3)n superlattices, such as flux-closure, vortice, and skyrmion. In this work, we employ deep potential molecular dynamics (MD) to investigate the dynamical response of the (PbTiO3)10/(SrTiO3)10 superlattice, supporting polar vortex arrays, to temperature and electric field at the atomic level. Our simulations reveal a unique phase transition sequence from ferroelectric-like to antiferroelectric-like to paraelectric in the in-plane direction as temperature increases. In the ferroelectric-like state, we observe field-driven reversible switching of in-plane polarization coupled with out-of-plane movements of vortex cores during MD simulations. In the antiferroelectric-like region, the polarization-electric field hysteresis loop exhibits a superparaelectric feature, showing nearly no loss. This behavior is attributed to a strong recovering force to form polar vortex arrays, dictated by the electrical and mechanical boundary conditions within the superlattice. The (PbTiO3)10/(SrTiO3)10 superlattice in the antiferroelectric-like state also demonstrates large in-plane susceptibility and tunability. The effect of Pb doping in the SrTiO3 layer on the topological structural transition in the superlattice is investigated. The weakened depolarization field in the PbTiO3 layers leads to new dipolar configurations, such as enlarged skyrmion bubble within $c$ domains in (PbTiO3)10/(Pb0.4Sr0.6TiO3)10, and we quantify their thermal and electrical responses through MD simulations. These quantitative atomistic insights will be useful for the controlled optimization of perovskite superlattices for various device applications.
@@ I Introduction
Ferroelectric materials, known for their spontaneous polarization that can be reversed by an external electric field, have been a focus of research due to their unique properties and diverse applications. Ferroelectrics have been utilized in non-volatile memory devices [1, 2], sensors [3], and actuators [4], capitalizing on their ability to retain polarization states in the absence of external bias and their strong coupling of the polarization to external stimuli. Over the past decade, considerable progress has been made in understanding and controlling the spontaneously formed topological polar structures in ferroelectric oxide heterostructures, including flux-closure [5],vortex [6], skyrmion [7], meron [8], and dipole wave [9, 10]. These nontrivial ferroelectric topological structures, the electric counterparts of their spin analogues, are formed due to the delicate competitions between electric, gradient and elastic energy terms.
Previously developed theoretical tools, including second-principles calculations and phase-field simulations, have achieved tremendous success in understanding and designing ferroelectric topological structures [6, 11, 7, 8, 9, 10, 12, 13, 14]. The main methods to control the dipolar topology include strain engineering and periodicity modulation. Taking (PbTiO3)n/(SrTiO3)n [(PTO)n/(STO)n] superlattice as an example, in-plane strain arising from the lattice mismatch between the superlattice and the substrate can influence the types of topological structures that emerge. For instance, vortex-type topological structures form under tensile strain [6], while dipolar skyrmions appear under slightly compressive strain [7]. The periodicity of the superlattice, defined by the number of PTO and STO layers, also has a significant impact on the formation of topological structures. For example, Hong et al. revealed through phase-field simulations that in (PTO)n/(STO)n, as $n$ increases, the domain structure evolves from $a_{1}/a_{2}$ to vortex and eventually to flux-closure [12]. However, the dynamical response of these polar topological structures to external electric fields at finite temperatures has rarely been explored at the atomic level. Furthermore, the impact of doping on the formation, stability, and dynamical properties of topological structures in ferroelectric superlattices remains largely unexplored.
@@ II.1 Deep potential of PSTO
The essence of the deep potential (DP) model is to construct a deep neural network that maps local atomic environments to atomic energies and forces, as extensively discussed in previous studies [20, 21, 22]. To train our DP model for PSTO solid solutions, we built a dataset containing various configurations of PTO, STO, and PbαSr1-αTiO3. The dataset includes 13,021 PTO structures in the $P4mm$ and $Pm\bar{3}m$ phases, 3,538 STO structures in the $I4/mcm$ and $Pm\bar{3}m$ phases, and 2,560 PSTO structures with compositions of $\alpha=0.25$, 0.5, and 0.75. This DP model accurately describes the phase diagram of PSTO across the entire temperature and composition range, and successfully reproduce the experimentally observed strain-induced topological structure transition [23]. We have developed an online notebook [24] on Github, which provides access to the training database, force field model, and training metadata.
@@ II.2 MD simulations
To simulate the (PTO)10/(STO)10 superlattice, we construct a 40$\times$20$\times$20 supercell containing 80,000 atoms, and set the in-plane lattice parameters of the supercell to $a_{\mathrm{IP}}=3.937$  and $b_{\mathrm{IP}}=3.929$. The effective strain is calculated as $\delta_{\mathrm{ave}}=(a_{\mathrm{IP}}+b_{\mathrm{IP}})/2a_{\mathrm{STO}}=0.60\%$, where $a_{\mathrm{STO}}=3.910$ is the in-plane ground-state lattice parameter of STO at 300 K in our MD simulations [23]. Under the condition of $\delta_{\mathrm{ave}}=0.60\%$, we observe the formation of polar vortex arrays formed in the PTO layers, consistent with experiments [11]. To simulate the (PTO)10/(PSTO)10 superlattice, Pb atoms are randomly doped into the STO layers. MD simulations in the constant-temperature constant-pressure ($NPT$) ensemble are carried out using the LAMMPS [25] package. In all simulations, a timestep of 2 fs is used for integrating the equations of motion. The system temperature is controlled using the Noé-Hoover thermostat, while the Parrinello-Rahman barostat is employed to maintain the pressure. During the simulation of temperature-driven phase transitions, the pressure is held at 1.0 bar. For each temperature, the system is first equilibrated for 50 ps, followed by a 50 ps production run. We find that this simulation duration is adequate to achieve convergence of the structural parameters.
We utilize the “force method” [26, 27] in classical MD to simulate the effects of external electric fields on the dynamical response of topological structures. In this method, the field-induced force $F_{i}$ on ion $i$ is the product of its Born effective charge (BEC) tensor $Z_{i}^{*}$ and the electric field strength $E$. To obtain the BEC tensors, we employ density functional perturbation theory [28, 29] to calculate the values for ions in the unit cells of the PTO tetragonal phase and the STO cubic phase, as reported in Table 1. The polarization of a unit cell at time $t$ can be expressed as:
@@ III.1 Temperature-driven phase transitions in superlattices
We examine the microscopic domain structures at several representative temperatures using MD simulations. At a low temperature of 100 K, the vortex cores are situated near the PTO/STO interfaces, forming a zigzag chain. The arrangement of alternating $+P_{z}$ and $-P_{z}$ domains, each approximately 5 unit cells wide, results in a cancellation of out-of-plane polarization. Interestingly, the zigzag vortex cores break global in-plane inversion symmetry, leading to alternating $+P_{x}$ and $-P_{x}$ domains of unequal size. This asymmetry prevents the polarization of $+P_{x}$ and $-P_{x}$ domains from completely canceling out, giving rise to a ferroelectric-like (FE) state with net polarization along the $x$ axis (Fig. 1c). As the temperature increases, we observe a collective shift of vortex cores away from the interfaces (Fig. 1b). At 300 K as shown in Fig. 1c, vortex cores are aligned along the $x$ axis, leading to equal volumes of $P_{x}$ and $-P_{x}$ domains. This configuration, termed antiferroelectric-like (AFE), features zero net in-plane polarization. The AFE state remains stable up to 550 K (Fig. 1f-g). At a higher temperature of 600 K, the entire superlattice transitions to a paraelectric state, characterized by randomly oriented local dipoles (Fig. 1h). The conterintuitive temperature-driven FE-AFE transition in the (PTO)10/(STO)10 superlattice is unusual and distinct from the behavior of single-crystal ferroelectrics, which rarely evolve into an antiferroelectric state with increasing temperature.
@@ III.2 Temperature-dependent polarization-electric field hysteresis loops
In contrast, the simulated $P_{x}$-$E_{x}$ loop at 300 K shows no remnant polarization and is nearly hysteresis-free (Fig. 3a). The S-shaped curve differs from the double hysteresis loop typically associated with antiferroelectric materials [33, 34]. Instead, it bears a resemblance to the behavior seen in “superparaelectric” systems [35, 36]. As illustrated in Fig. 3b, the initial configuration features $+P_{x}$ and $-P_{x}$ domains of equal volume with aligned vortex cores. The application of $+E_{x}$ drives opposing movements of neighboring vortex cores. This process is coupled with the growth of $P_{x}$ domains (Fig. 3c). At a sufficiently high field of 0.4 MV/cm, the vortex state evolves smoothly to a dipole wave state (Fig. 3d) and eventually to a single-domain state with $+P_{x}$. As the field is reduced, the superlattice undergoes successive transitions from an $a$-domain to dipole waves and back to vortex arrays in the PTO layers (Fig. 3f). At zero field, polar vortex arrays return to the original state without net in-plane polarization (Fig. 3g). The spontaneous formation of polar vortex lattices indicates a one-to-one correspondence between the topological structure and the mechanical and electrical boundary conditions, which include factors such as epitaxial strain and lattice periodicity, within a superlattice at a given temperature. These conditions collectively determine the energetically favorable polar configuration and act as a restorative force, reforming vortex arrays upon field removal. This explains the hysteresis-free, superparaelectric-like $P_{x}$-$E_{x}$ curve.
@@ III.3 In-plane dielectric susceptibility and tunability
We extract the temperature dependence of the in-plane dielectric constant $\epsilon_{\mathrm{eff}}$ from the $P_{x}$-$E_{x}$ loops at various temperatures. For a typical ferroelectric, the dielectric constant often undergoes a discontinuous change at the FE-PE phase transition [37]. Here, we find that the in-plane dielectric constant of the superlattice also exhibits a discontinuous change at the FE-AFE transition (Fig. 4a). In contrast, the dielectric constant remains continuous at the AFE-PE transition, where no global symmetry breaking occurs. Previous studies have demonstrated that the position of the vortex core can be modulated by strain [10, 38], suggesting that the FE-AFE transition temperature (or the peak of the dielectric constant) can be tuned to accommodate different application requirements.
Furthermore, we compare the in-plane dielectric constant of the superlattice and the out-of-plane dielectric constant of PTO under varying electric fields at 300 K. As plotted in Fig. 4b, our results indicate that the dielectric constant of the superlattice is highly tunable, reaching a maximum value of 3900 in units of $\epsilon_{0}$ (vacuum permittivity), while the dielectric constant of PTO is only 170, consistent with experimental values reported in the literature [39, 40]. This significant enhancement in the in-plane dielectric constant of the superlattice, compared to bulk PTO, underscores the potential of these topological structures for applications requiring high dielectric susceptibility and tunability.

### v4

@@ Abstract
Many dipolar topological structures have been experimentally demonstrated in (PbTiO3)n/(SrTiO3)n superlattices, such as flux-closure, vortice, and skyrmion. In this work, we employ deep potential molecular dynamics (MD) to investigate the atomic-level dynamical response of the (PbTiO3)10/(SrTiO3)10 superlattice, which hosts polar vortex arrays, to variations in temperature and electric field. Our simulations reveal a unique phase transition sequence from ferroelectric-like to antiferroelectric-like to paraelectric in the in-plane direction as temperature increases. In the ferroelectric-like state, we observe field-driven reversible switching of in-plane polarization coupled with out-of-plane movements of vortex cores. In the antiferroelectric-like region, the polarization-electric field hysteresis loop exhibits a superparaelectric feature, showing nearly no loss. This behavior is attributed to a strong recovering force to form polar vortex arrays, dictated by the electrical and mechanical boundary conditions within the superlattice. The (PbTiO3)10/(SrTiO3)10 superlattice in the antiferroelectric-like state also demonstrates large in-plane susceptibility and tunability. The effect of Pb doping in the SrTiO3 layer on the topological structural transition in the superlattice is investigated. The weakened depolarization field in the PbTiO3 layers leads to new dipolar configurations, such as enlarged skyrmion bubble within $c$ domains in (PbTiO3)10/(Pb0.4Sr0.6TiO3)10, and we quantify their thermal and electrical responses using MD simulations. These quantitative atomistic insights will be useful for the controlled optimization of perovskite superlattices for various device applications.
@@ I Introduction
Ferroelectric materials, known for their spontaneous polarization that can be reversed by an external electric field, have been a focus of research due to their unique properties and diverse applications. Ferroelectrics have been utilized in non-volatile memory devices [1, 2], sensors [3], and actuators [4], capitalizing on their ability to retain polarization states in the absence of external bias and their strong coupling of the polarization to external stimuli. Over the past decade, considerable progress has been made in understanding and controlling the spontaneously formed topological polar structures in ferroelectric oxide heterostructures, including flux-closure [5],vortex [6], skyrmion [7], meron [8], and dipole wave [9, 10]. These nontrivial ferroelectric topological structures, often regarded as the electric counterparts of their spin analogues, are formed due to the delicate competitions between electric, gradient and elastic energy terms.
Previously, developed theoretical tools, such as second-principles calculations and phase-field simulations, have been highly successful in advancing the understanding and designing ferroelectric topological structures [6, 11, 7, 8, 9, 10, 12, 13, 14, 15, 16, 17]. More recently, Yuan et al.  revealed the existence of a field-induced hexagonal-lattice skyrmion crystal in PbTiO3 thin films through phase-field simulations [18]. Aramberri et al.  employed a second-principles approach to investigate the dynamics of electric skyrmions, demonstrating the quasiparticle nature of these electric bubbles [19]. The main methods to control the dipolar topology include strain engineering and periodicity modulation. Taking (PbTiO3)n/(SrTiO3)n [(PTO)n/(STO)n] superlattice as an example, in-plane strain arising from the lattice mismatch between the superlattice and the substrate can influence the types of topological structures that emerge. For instance, vortex-type topological structures form under tensile strain [6], while dipolar skyrmions appear under slightly compressive strain [7]. The periodicity of the superlattice, defined by the number of PTO and STO layers, also has a significant impact on the formation of topological structures. For example, Hong et al. . revealed through phase-field simulations that in (PTO)n/(STO)n, as $n$ increases, the domain structure evolves from $a_{1}/a_{2}$ to vortex and eventually to flux-closure [12]. However, the impact of doping on the formation, stability, and dynamical properties of topological structures in ferroelectric superlattices remains largely unexplored at the atomic level.
@@ II.1 Deep potential of PSTO
The essence of the deep potential (DP) model is to construct a deep neural network that maps local atomic environments to atomic energies and forces, as extensively discussed in previous studies [25, 26, 27]. To train our DP model for PSTO solid solutions, we built a dataset containing the DFT energies and forces of various configurations of PTO, STO, and PbαSr1-αTiO3. The dataset includes 13,021 PTO structures in the $P4mm$ and $Pm\bar{3}m$ phases, 3,538 STO structures in the $I4/mcm$ and $Pm\bar{3}m$ phases, and 2,560 PSTO structures with compositions of $\alpha=0.25$, 0.5, and 0.75. All DFT calculations are performed with the Vienna Ab initio Simulation (VASP) package [28, 29] using the projected augmented wave (PAW) method [30, 31]. Perdew-Burke-Ernzerhof functional modified for solids (PBEsol) [32] within the generalized gradient approximation is chosen as the exchange-correlation functional. An energy cutoff of 800 eV and a $k$-point spacing of 0.3 Å-1 are sufficient to converge the energy and atomic forces. Additional details on model development can be found in ref. [33]. Density functional perturbation theory (DFPT) is employed to directly calculate the first-order response of the wavefunction (charge density) to an applied electric field for computing the Born effective charges [34, 35]. The energy cutoff and $k$-point spacing settings in the DFPT calculations are kept the same as those used in constructing the training dataset. This DP model accurately describes the phase diagram of PSTO across the entire temperature and composition range, and successfully reproduce the experimentally observed strain-induced topological structure transition [33]. We have developed an online notebook [36] on Github, which provides access to the training database, force field model, and training metadata.
@@ II.2 MD simulations
To simulate the (PTO)10/(STO)10 superlattice, we construct a 40$\times$20$\times$20 supercell containing 80,000 atoms. We note that PBEsol is chosen as the exchange-correlation functional for constructing the DFT database used to train the DP model. Since the ground-state lattice constants of typical perovskite oxides, such as PbTiO3 and SrTiO3, predicted by PBEsol deviate slightly from experimental values, it is expected that the theoretical mechanical boundary conditions leading to the formation of a specific topological structure will differ slightly from those observed experimentally. In this study, we focus specifically on the pure vortex state. We explored various combinations of in-plane lattice parameters, $a_{\rm IP}$ and $b_{\rm IP}$, and found that $a_{\mathrm{IP}}=3.937$ and $b_{\mathrm{IP}}=3.929$ leads to well-defined polar vortex arrays. In reference to the in-plane ground-state lattice parameter of PTO at 300 K ($a_{\mathrm{PTO}}=3.919$) in our MD simulations, the strain state is $\delta_{x}=0.46\%$ and $\delta_{y}=0.26\%$, and the average strain is $\delta_{\mathrm{ave}}=0.36\%$. The value of $\delta_{\mathrm{ave}}$ agrees reasonably well with the experimental value of 0.3% that yields a pure vortex state [37] and the 1.1% strain that leads to a mixed state of $a_{1}$/$a_{2}$ and vortex arrays [6, 38]. To simulate the (PTO)10/(PSTO)10 superlattice, Pb atoms are randomly doped into the STO layers. MD simulations in the constant-temperature constant-pressure ($NPT$) ensemble are carried out using the LAMMPS package [39]. We note that LAMMPS allows for independent specification of all six components of an external stress tensor. This capability enables us to control each of the six dimensions of the simulation box independently. In our simulations, we fixed the in-plane lattice constants while allowing the out-of-plane lattice constant to adjust in response to the target pressure. In all simulations, a timestep of 2 fs is used for integrating the equations of motion. The system temperature is controlled using the Noé-Hoover thermostat, while the Parrinello-Rahman barostat is employed to maintain the pressure. During the simulation of temperature-driven phase transitions, the pressure is held at 1.0 bar. For each temperature, the system is first equilibrated for 50 ps, followed by a 50 ps production run. We find that this simulation duration is adequate to achieve convergence of the structural parameters.
We utilize the “force method” [40, 41] in classical MD to simulate the effects of external electric fields on the dynamical response of topological structures. In this method, the field-induced force $F_{i}$ on ion $i$ is the product of its Born effective charge (BEC) tensor $Z_{i}^{*}$ and the electric field strength $E$. To obtain the BEC tensors, we employ the DFPT method [34, 35] to calculate the values for ions in the unit cells of the PTO tetragonal phase and the STO cubic phase, as reported in Table 1. The polarization of a unit cell at time $t$ can be expressed as:
@@ III.1 Temperature-driven phase transitions in superlattices
We examine the microscopic domain structures at several representative temperatures using MD simulations. At a low temperature of 100 K, the vortex cores are situated near the PTO/STO interfaces, forming a zigzag chain. The arrangement of alternating $+P_{z}$ and $-P_{z}$ domains, each approximately 5 unit cells wide, results in a cancellation of out-of-plane polarization. Interestingly, the zigzag vortex cores break global in-plane inversion symmetry, leading to alternating $+P_{x}$ and $-P_{x}$ domains of unequal size. This asymmetry prevents the polarization of $+P_{x}$ and $-P_{x}$ domains from completely canceling out, giving rise to a ferroelectric-like (FE) state with net polarization along the $x$ axis (Fig. 1c). We note that it is more rigorous to label this domain structure featuring unequal volume fractions of $+P_{x}$ and $-P_{x}$ domains as “ferrielectric-like”. For convenience, we will continue to use the term “ferroelectric-like” in the following discussion. As the temperature increases, we observe a collective shift of vortex cores away from the interfaces (Fig. 1d). At 300 K as shown in Fig. 1e, vortex cores are aligned along the $x$ axis, leading to equal volumes of $P_{x}$ and $-P_{x}$ domains. This configuration, termed antiferroelectric-like (AFE), features zero net in-plane polarization. The AFE state remains stable up to 550 K (Fig. 1f-g). At a higher temperature of 600 K, the entire superlattice transitions to a paraelectric state, characterized by randomly oriented local dipoles (Fig. 1h). The conterintuitive temperature-driven FE-AFE transition in the (PTO)10/(STO)10 superlattice is unusual and distinct from the behavior of single-crystal ferroelectrics, which rarely evolve into an antiferroelectric state with increasing temperature. Additionally, a careful examination of the polar structures around the AFE-PE transition temperature reveals the presence of an intermediate topological phase. As shown in Fig. 2, the transition is not sharp, mediated by a complex polar liquid phase characterized by the loss of long-range translational order along the $y$ axis. This observation aligns with the recently reported two-step melting process, which involves an intermediate hexatic-like phase [42, 43].
The collective movement of vortex cores is a manifestation of the vortexon mode [38, 44]. We track the position of the vortex core in MD simulations by identifying the extremal points of vorticity, as illustrated in Fig. 3a. We introduced a new order parameter $\xi$, defined as the out-of-plane distance between the vortex core to the center of the PbTiO3 layer. Figure  3b presents the temperature dependence of the ensemble-averaged value of $\xi$ from MD simulations, revealing a displacive, Landau-type first-order phase transition. Furthermore, we perform a Fourier transform on the time evolution of $\xi$ at 300 K (Fig. 3c-d) and obtain the frequency spectrum, which shows a peak at 0.1 THz, closely aligning with the experimental value of 0.08 THz [38]. These results confirm the presence of vortexon modes and validate the accuracy of the DP model utilized in this study. Moreover, the displacement of the vortex core affects the ratio of $+P_{x}$ domains and $-P_{x}$ domains, thereby influencing the magnitude of the total $P_{x}$. Additionally, the (local) polarization of PbTiO3 also scales with $\sqrt{T_{c}-T}$, where $T_{c}$ is the intrinsic Curie temperature. The superposition of these two effects make $P_{x}$ decreases more rapidly, as shown in Fig. 1b.
@@ III.2 Temperature-dependent polarization-electric field hysteresis loops
In contrast, the simulated $P_{x}$-$E_{x}$ loop at 300 K shows no remnant polarization and is nearly hysteresis-free (Fig. 5a). The S-shaped curve differs from the double hysteresis loop typically associated with antiferroelectric materials [47, 48]. Instead, it bears a resemblance to the behavior seen in “superparaelectric” systems [49, 50]. Previously, Aramberri et al.  employed a second-principles method to simulate (PTO)4/(STO)4 superlattices and also observed the evolution from a double hysteresis loop at 0 K to a nonhysteric S-shaped curve at room temperature [51]. As illustrated in Fig. 5b, the initial configuration features $+P_{x}$ and $-P_{x}$ domains of equal volume with aligned vortex cores. The application of $+E_{x}$ drives opposing movements of neighboring vortex cores. This process is coupled with the growth of $P_{x}$ domains (Fig. 5c). At a sufficiently high field of 0.4 MV/cm, the vortex state evolves smoothly to a dipole wave state (Fig. 5d) and eventually to a single-domain state with $+P_{x}$. As the field is reduced, the superlattice undergoes successive transitions from an $a$-domain to dipole waves and back to vortex arrays in the PTO layers (Fig. 5f). At zero field, polar vortex arrays return to the original state without net in-plane polarization (Fig. 5g). The spontaneous formation of polar vortex lattices indicates a one-to-one correspondence between the topological structure and the mechanical and electrical boundary conditions, which include factors such as epitaxial strain and lattice periodicity, within a superlattice at a given temperature. These conditions collectively determine the energetically favorable polar configuration and act as a restorative force, reforming vortex arrays upon field removal. This explains the hysteresis-free, superparaelectric-like $P_{x}$-$E_{x}$ curve.
@@ III.3 In-plane dielectric susceptibility and tunability
We extract the temperature dependence of the in-plane static dielectric constant $\epsilon_{\mathrm{eff}}$ from the $P_{x}$-$E_{x}$ loops at various temperatures. For a typical ferroelectric, the dielectric constant often undergoes a discontinuous change at the FE-PE phase transition [52]. Here, we find that the in-plane dielectric constant of the superlattice also exhibits a discontinuous change at the FE-AFE transition (Fig. 6a). In contrast, the dielectric constant remains nearly continuous at the AFE-PE transition, likely due to the intermediate polar liquid phase (Fig. 2) that smooths the dielectric response across the transition temperature. Previous studies have demonstrated that the position of the vortex core can be modulated by strain [10, 44], suggesting that the FE-AFE transition temperature (or the peak of the dielectric constant) can be tuned to accommodate different application requirements.
We compare the in-plane dielectric constant of the superlattice and the out-of-plane dielectric constant of PTO under varying electric fields at 300 K. For a given electric field, we first obtained the equilibrium domain structure with finite-field MD simulations. Subsequently, we applied a much smaller in-plane probing electric field of approximately 0.02 MV/cm (in addition to the background field) to perturb the polarization. The in-plane dielectric constant is then computed from the slope of the polarization-electric field curve. The probing electric field for susceptibility measurements is significantly smaller than the background electric field to ensure the validity of linear response theory. As plotted in Fig. 6b, our results indicate that the dielectric constant of the superlattice is highly tunable, reaching a maximum value of 3900 in units of $\epsilon_{0}$ (vacuum permittivity), while the dielectric constant of PTO is only 170, consistent with experimental values reported in the literature [53, 54]. This significant enhancement in the in-plane dielectric constant of the superlattice, compared to bulk PTO, underscores the potential of these topological structures for applications requiring high dielectric susceptibility and tunability.